OpenStudy (anonymous):

14 is what percent of 28?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

how did you come up with that?

OpenStudy (anonymous):

14/28 = .5 * 100 = 50%

OpenStudy (anonymous):

14 is half of 28, therefore 50%

OpenStudy (anonymous):

If you read that statement as an equation you have 14=X*28, idk if you are in that kind of math yet thoug

OpenStudy (migitmack):

idk i doubt it

OpenStudy (anonymous):

I really appreciate the help, im going to the 8 th grade this yr n im taking some classes to try n improve my math skills...so thnk u

OpenStudy (anonymous):

good for you in preparing.

OpenStudy (migitmack):

how comes 50% would be 25 ah nvm